OpenStudy (anonymous):

[(x+4)-2y] squared im stuck !

OpenStudy (anonymous):

\[[(x+4)-2y]^2\]Is this the question? just making sure

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[(x+4-2y)^2=(x+4-2y)(x+4-2y)\]

OpenStudy (anonymous):

you have to go one term at a time.

OpenStudy (anonymous):

Almost like you are using FOIL but with 3 terms??

OpenStudy (anonymous):

\[(a+b-c)(a+b-c)=aa+ab-ac+ba+bb-bc-ca-cb+cc\]

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

same principle

OpenStudy (anonymous):

I broke apart the ( ) but you could foil it and treat x+4 as 1 term

OpenStudy (anonymous):

I find it easier to make a mistake that way

OpenStudy (anonymous):

oh ok think i got it now, Thank you so much

OpenStudy (anonymous):

yup